### v1

@@ Discussion
For the $m$-th image-text pair, a patch-concept similarity matrix $\mathbf{A}\in\mathbb{R}^{r\times w}$, where $w$ is the number of concepts in the $m$-th image-text pair. Then, for the $i$-th patch and the $j$-th concept, the alignment score is:
$(\mathbf{a})_{ij}=\frac{(\mathbf{g}_{m,j})^{T}\mathbf{t}_{m,i}}{||\mathbf{g}_{m,j}||\cdot||\mathbf{t}_{m,i}||}$ (7)
$(\mathbf{A})_{ij}=\text{log}\frac{1}{1+e^{z^{m,n}(-t_{g}\mathbf{a}_{i,j}+b_{g})}}$ (8)
where $z^{m,n}=1$ if the $m$-th image and the $n$-th text are paired. Otherwise, $z^{m,n}=-1$. $t_{g}$ and $b_{g}$ are the logit scale and logit bias for the patch-concept alignment.
For each image $I_{m}$ and text $T_{n}$, the similarity score is:
$S(I_{m},T_{n})=\frac{1}{w}\sum_{j=1}^{w}\max_{i}(\mathbf{A}_{ij})$ (9)
@@
Thus, the PC-Align loss in a mini-batch of $|\mathcal{B}|$ image-text-concept triples is:
$\mathcal{L}_{\text{PC-Align}}=-\frac{1}{|\mathcal{B}|}\sum_{m=1}^{|\mathcal{B}|}\sum_{n=1}^{|\mathcal{B}|}S(I_{m},T_{n})$ (10)
@@
$\mathcal{L}=\mathcal{L}_{\text{IT-Align}}+\alpha\mathcal{L}_{PC-Align}$ (11)
where $\alpha$ is a hyper-parameter to decide the weight of the PC-Align loss during the pre-training.
@@
In this section, we elucidate the inference strategy employed in our model, which integrates both local and global information to enhance the zero-shot classification performance [37]. This approach is particularly advantageous in the medical domain, where nuanced and context-specific details often play a crucial role in accurate diagnosis.
The integration of local and global information is inspired by the need to capture both the holistic context of an image and the specific details that may be critical for identifying particular medical conditions. While global image-caption alignment provides a broad understanding of the visual and textual data, local patch-concept alignment allows for a more detailed examination of specific regions within an image that correspond to medical concepts.
Our inference strategy employs a dual-classifier system consisting of a global classifier and a local classifier. The global classifier leverages the overall image features, while the local classifier focuses on the alignment between image patches and text-derived concepts. By combining these classifiers, our model can make more informed decisions that account for both general and specific information.
During inference, the model processes input images through an image encoder to extract image features, denoted as $\mathbf{i}_{m}$. Simultaneously, the text encoder processes the associated concepts to derive text features, $\mathbf{t}_{m}$. The global classifier computes logits based on the similarity between the normalized global image features and the global class representations obtained during pre-training. For the local classifier, the model evaluates patch-level features against concept representations. This involves computing a patch-concept similarity matrix, which quantifies the alignment between image patches and text-derived concepts. The top-k alignment scores are aggregated to form the local logits, reflecting the degree of correspondence between image patches and medical concepts.
The final classification decision is derived by combining the outputs of the global and local classifiers. A weighted sum of the global and local logits is computed, controlled by a hyper-parameter $\beta$, which balances the contribution of each classifier:
$=\beta\cdot\text{local-logits}+(1-\beta)\cdot\text{global-logits}$-logits (12)
This integration allows the model to harness the strengths of both global and fine-grained local information, thereby improving the robustness and accuracy of zero-shot classification in medical imaging tasks.
@@
In this section, we first describe the dataset used to pre-train the medical vision-language model. Based on the pre-trained model, we conduct extensive experiments on downstream tasks. Besides, we describe the baseline medical vision-language models in detail. Finally, we cover the implementation details in pre-training the proposed ConceptCLIP.
@@
In our pre-training process, the proposed MedConcept-23M is used. In detail, it contains 23 million image-text-concept triples, where the image-text pairs are collected from PMC-OA and concepts are aligned to the UMLS concepts and extracted from the text.
@@
We assess the zero-shot, linear probe (using 1%, 10%, and 100% of the training data), and full fine-tuning capabilities of ConceptCLIP and previous state-of-the-art medical CLIP models across 36 datasets spanning 10 image modalities. In the zero-shot setting, no ground truth labels are provided. We follow the zero-shot protocol of the CLIP model [4], where the text encoder and image encoder are used to encode the medical images and labels (using a prompt template to form a sentence), calculate the similarity, and assign labels to the images. For linear probes, we first extract features from each image using a specific image encoder from a pre-trained model and employ these image features and their labels to perform logistic regression. The trained logistic regression models are subsequently used to test model performance. This experiment aims to evaluate each model’s ability to extract image features. In the full fine-tuning setting, each model is concatenated with a classification head and fine-tuned on a specific dataset with all parameters adjustable. We use the AUC score for model evaluation to ensure that the performance is not influenced by thresholds.
Cross-modal image-text retrieval
@@
Cross-modal image-text retrieval includes tasks such as medical image-to-text and text-to-image retrieval. To evaluate the performance of ConceptCLIP across various medical image modalities, we utilize two datasets: PMC-9K and Quilt-1M [16], to assess the models’ cross-modal retrieval capabilities in a mix of image modalities and pathology images. We use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to evaluate each model’s performance.
Medical report generation
@@
Medical report generation aims to alleviate the workload of doctors by automatically generating a report given a medical image. In our study, following the shallow alignment in the R2GenGPT framework [38], we replace the vision encoder with that of each medical CLIP model and use the same training scheme to fine-tune the report generation model. We conduct these experiments on the MIMIC-CXR [39] and IU X-Ray [40] datasets. To comprehensively evaluate each model’s performance, we use natural language generation metrics: BLEU-1,2,3,4, METEOR, ROUGE-L, CIDEr, and clinical efficacy metrics: Micro Precision, Micro Recall, and Micro F1.
Medical visual question answering
@@
Medical Visual Question Answering (VQA) serves as a bridge between AI systems and humans. Given a medical image, this task involves providing a question about the image to the system, which should then answer the question based on the image. In our experiments, we use the METER framework [41] to evaluate different models on the VQA task, employing the SLAKE [42] and VQA-RAD [43] datasets. Following METER’s evaluation metrics, we use Closed Accuracy, Open Accuracy, and Overall Accuracy to assess each model’s performance.
Whole-slide image analysis
@@
To explore our model’s ability to understand whole-slide images, we conduct experiments on 6 related datasets for 9 subtasks in Cancer Diagnosis, Molecular Subtyping, and Survival Prediction, each of which can be regarded as a classification task, with a whole-slide image as input to the vision encoder and a predicted label as output. Given the high resolution of whole-slide images, we follow the conventional approach by segmenting the whole-slide image into smaller patches and using each pre-trained vision encoder to extract features from each image patch. These features are input into the ABMIL model [44] for multiple instance learning (MIL). The model outputs a predicted label. Similar to the Medical Image Diagnosis task, we use the AUC score to evaluate different models in Cancer Diagnosis and Molecular Subtyping, and concordance index (C-index) for survival analysis.
Medical concept annotation
@@
We evaluate to what extent can the proposed ConceptCLIP annotate fine-grained concepts for medical images, hence facilitating the development of concept-based explainable artificial intelligence. The performance of medical concept annotation is evaluated using the concept labels of the datasets. For each concept, two positive and negative prompts are designed to calculate the similarity with input images to get the predicted output. Specifically, we adopt five datasets with clinical concept annotations to evaluate the zero-shot concept annotation performance, including dermoscopic images, blood cell images (hematology), breast ultrasound, and chest CT.
Derm7pt [10]: A dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions in dermatology. The test set is adopted for evaluation. Only the dermoscopic images are considered in this paper for concept annotations. The seven clinical concepts include “Pigment network”, “Dots and globules”, “Pigmentation”, “Streaks”, “Regression structures”, “Blue-whitish veil”, and “Vascular structures”.
SkinCon [9]: A skin disease dataset densely annotated by experts for fine-grained model debugging and analysis. We choose 22 concepts that have at least 50 images representing the concept from the Fitzpatrick 17k (F17k) [45] subset. 350 images are sampled for evaluation. The clinical concepts include “Papule”, “Plaque”, “Pustule”, “Bulla”, “Patch”, “Nodule”, “Ulcer”, “Crust”, “Erosion”, “Atrophy”, “Exudate”, “Telangiectasia”, “Scale”, “Scar”, “Friable”, “Dome-shaped”, “Brown (Hyperpigmentation)”, “White (Hypopigmentation)”, “Purple”, “Yellow”, “Black”, and “Erythema”.
WBCAtt [11]: This dataset is a densely annotated dataset for white blood cell (WBC) recognition, containing 11 morphological attributes for 10,298 cell images. The test set which contains 3,099 images is used for evaluation. There are 11 concepts, including “Cell size”,“Cell shape”,“Nucleus shape”,“Nuclear cytoplasmic ratio”,“Chromatin density”, “Cytoplasm vacuole”, “Cytoplasm texture”, “Cytoplasm colour”, “Granule type”, “Granule colour”, and “Granularity”.
BrEaST [12]: A breast ultrasound dataset containing 256 images with fine-grained concepts for breast lesion analysis. Seven BI-RADS descriptors are adopted as the evaluated concepts, including “Shape”, “Magrin”, “Echogenicity”, “Posterior features”, “Halo”, “Calcifications”, and “Skin thickening”.
LUNA16 [13]: A curated version of LIDC-IDRI [46] dataset, which contains lung images with clinical concepts. 1,185 slices of lesions from the chest CTs with six clinical concepts are used for evaluation. The concepts include “Calcification”, “Lobulation”, “Margin”, “Sphericity”, “Spiculation”, and “Texture”.
Concept-based data analysis
To conduct concept-based data analysis, concept presence difference is calculated for datasets of four modalities, including NLM-TB (chest X-ray), CBIS-DDSM (mammography), UBIBC (breast ultrasound), and MedFMC-Colon (pathology). Specifically, for each dataset, assume there are two image sets of different classes (e.g., tuberculosis and normal in NLM-TB dataset), denoted as $I_{+}=\{I_{1+},I_{2+},...,I_{M+}\}$ and $I_{-}=\{I_{1-},I_{2-},...,I_{N-}\}$, and the pre-defined concept list $C$ (e.g., cavitation, consolidation, etc.), where $M$ and $N$ are the numbers of images in the positive and negative image sets, respectively, and $C$ are generated by a large language model (i.e., GPT-4 [47]) since the dataset does not have concept labels. We adopt the proposed ConceptCLIP to annotate the presence of the concepts for each given image. The concept presence proportion of each image set is computed by $\frac{N_{c_{i}}}{M}$ and $\frac{N_{c_{i}}}{N}$, respectively, here $N_{c_{i}}$ is the number of images where a specific concept $c_{i}$ is present. Then the concept presence difference of concept $c_{i}$ is defined as the difference between the two concept presence proportions, i.e., $D_{c_{i}}=\frac{N_{c_{i}}}{M}-\frac{N_{c_{i}}}{N}$. A positive value of $D_{c_{i}}$ means concept $c_{i}$ is more likely to appear in the positive image set $I_{+}$, and vice versa.
Inherently interpretable model employment and evaluation
@@
Building inherently interpretable models is essential for deploying AI in medicine since it is more trustworthy for doctors and patients during clinical diagnosis. In this paper, we employ the Concept Bottleneck Model (CBM) [22] to build the inherently interpretable model based on our method. Inspired by previous works [23, 24], we use ConceptCLIP to calculate the cosine similarity between input images and medical concepts, then the similarity is used as the input to the concept bottleneck layer, which maps the concept similarity to the final prediction. The concept bottleneck layer is trained using the class labels. For concept-class association, the weights of each neuron within the linear layer can be regarded as the contribution to the final prediction. To evaluate the effectiveness of our method, we compare ConceptCLIP with other vision-language models of both the natural and medical domains on four datasets of different modalities, including dermoscopy, hematology, ultrasound, and CT.
Disease-level concept inspection
As shown in Fig. 5, pneumonia is selected as the inspected disease. For concept inspection on a single dataset, RSNA [48] dataset is adopted, which only includes classes of No finding and Pneumonia. For disease-level concept inspection, we collect datasets that also include the two classes, i.e., VinDr-CXR [49] and VinDr-PCXR [50], where VinDr-PCXR focus on the chest X-rays of children. It is noteworthy that VinDr-CXR and VinDr-PCXR contain other classes such as Lung tumor and Tuberculosis, but we only consider the images of No Finding and Pneumonia to analyze the pneumonia disease. For each dataset, we use ConceptCLIP to obtain the concept-disease association. Then we average the weights of the concept bottleneck layer across all datasets to derive the final top relevant concepts for pneumonia.
General domain models and the state-of-the-art medical models
CLIP [4] is a vision-language model pre-trained on large-scale image-text pairs via contrastive learning in the general domain, which showcases strong ability in zero-shot classification. We take the version of CLIP-ViT-B-16 pre-trained on LAION-2B [51] under the framework of OpenCLIP [52].
SigLIP [5] is a variant of the CLIP model, which is pre-trained via contrastive learning using the sigmoid loss. In our experiments, we use the version of ViT-SO400M-14-SigLIP (SigLIP-400M).
PMC-CLIP [7] is a vision-language model tailored for medicine. The authors collected PubMed open access articles and conducted a series of data processing operations, producing a dataset with 1.6 million medical image-text pairs. The dataset is used to pre-train a CLIP model, thus resulting in the PMC-CLIP model. We use the origin checkpoint [^2] https://huggingface.co/datasets/axiong/pmc_oa_beta/blob/main/checkpoint.pt that the authors provided.
BiomedCLIP [8] is a medical vision-language model pre-trained on PMC-15M, a dataset with 15 million medical image-text pairs sourced from PMC-OA. We take the released checkpoint [^3] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224 for evaluation.
Implementation details
We develop ConceptCLIP by utilizing SigLIP-ViT-400M-16444https://huggingface.co/timm/ViT-SO400M-14-SigLIP-384 as the visual encoder and PubMedBERT [^5] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as the text encoder. Each input image is resized to $336\times 336$ pixels. During pre-training, we employ the AdamW optimizer [53] with a learning rate of $5\times 10^{-4}$. The model is initially pre-trained on H800 GPUs with a batch size of 12,288 for 32 epochs on the MedConcept-23M dataset, without the PC-Align loss. Subsequently, we conduct further pre-training with the PC-Align loss on this dataset with a batch size of 6,144 and learning rate of $3\times 10^{-4}$ for 20 epochs. The weight $\alpha$ during the pre-training is set to 0.5.
In medical image diagnosis, for zero-shot classification, we set $\beta$ to 0.5 so that the prediction can consider the information from both global information and local information. For different datasets, we choose different “Top-K” values and prompts as described in LABEL:tab:zero_shot_settings. For linear probes, for all experiments, we use the LogisticRegression function provided in scikit-learn [^6] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, with inverse of regularization strength is 0.316, maximum iteration is 1,000, random state is 1, and all other parameters are default. For full fine-tuning experiments, for all experiments, the batch size is 64, epochs is 20, learning rate is 1e-4. We use AdamW optimizer for training the model.
In the visual question answering task, we use AdamW as the optimized. For SigLIP-400M and ConceptCLIP, the batch size is 32, learning rate is 1e-6. For other models, the batch size is 4, learning rate is 5e-6. All experiments have the same epoch number 50.
In whole-slide image analysis, the batch size is 1, learning rate is 2e-4, epochs are 30, and the cosine learning rate scheduler is used.
@@
AUC
The Area Under the Receiver Operating Characteristic Curve (AUC) is a performance measurement for classification problems at various threshold settings. It is widely used in medical image diagnosis and explainable AI tasks to evaluate the ability of a model to distinguish between classes.
Accurary
Accuracy (ACC) is the proportion of true results among the total number of cases examined. In the context of visual question answering tasks, we use closed accuracy, open-accuracy, and overall accuracy to assess the model’s ability to select the correct answer from a candidate answer set. The formula is:
$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$
C-Index
The concordance index (C-Index) measures the predictive accuracy of a survival model. It is particularly used in survival prediction tasks in whole-slide image analysis to evaluate how well the model predicts the order of events. It is calculated as the proportion of all usable patient pairs whose predictions and outcomes are concordant.
Recall
Recall, or sensitivity, measures the ability of a model to correctly identify all relevant instances. In cross-modal retrieval tasks, we use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to assess the model’s retrieval performance at different levels. The formula is:
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
BLEU
The Bilingual Evaluation Understudy (BLEU) score [54] is a metric for evaluating the quality of text generated by a machine, such as translations or reports. It is used for medical report generation to assess the accuracy and fluency of the generated text. The formula for BLEU is:
$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
where $p_{n}$ is the precision for n-grams, $w_{n}$ are weights, and BP is the brevity penalty.
METEOR
The Metric for Evaluation of Translation with Explicit ORdering (METEOR) [55] is another metric for evaluating text generation quality, focusing on precision, recall, and alignment of phrases. It is used for medical report generation to provide a more nuanced assessment of generated text quality.
ROUGE-L
The Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L) [56] measures the longest common subsequence (LCS) between the generated text and reference texts. It is used for evaluating medical report generation to capture the overlap in content. The formula for ROUGE-L is:

### v2

@@ Discussion
For the $m$-th image and the $n$-th text, we define a region-concept similarity matrix $\mathbf{A}\in\mathbb{R}^{r\times w}$, where $w$ represents the number of concepts in the $n$-th text. The alignment score $\mathbf{a}_{ij}$ for the $i$-th region in the $m$-th image and the $j$-th concept in the $n$-th text is calculated as:
$\mathbf{a}_{ij}=\frac{(\mathbf{g}_{n,j})^{T}\mathbf{i}_{m,i}}{||\mathbf{g}_{n,j}||\cdot||\mathbf{i}_{m,i}||}.$ (5)
The similarity matrix element $\mathbf{A}_{ij}$ is then expressed as:
$\mathbf{A}_{ij}=\text{log}\frac{1}{1+e^{-t_{l}\mathbf{a}_{i,j}+b_{l}}}.$ (6)
The parameters $t_{l}$ and $b_{l}$ are the logit scale and logit bias for the local region-concept alignment, respectively.
For the $m$-th image $I_{m}$ and the $n$-th text $T_{n}$, the similarity score is:
$S(I_{m},T_{n})=\frac{1}{w}\sum_{j=1}^{w}\max_{i}(\mathbf{A}_{ij}).$ (7)
@@
Thus, the RC-Align loss in a mini-batch of $|\mathcal{B}|$ image-text-concept triplets is:
$\mathcal{L}_{\text{RC-Align}}=-\frac{1}{|\mathcal{B}|}\sum_{m=1}^{|\mathcal{B}|}\sum_{n=1}^{|\mathcal{B}|}z^{m,n}\cdot S(I_{m},T_{n}).$ (8)
This ensures that the representations of the most similar image region and textual concept are closer when the image and the text are paired, and further apart otherwise.
@@
$\mathcal{L}=\mathcal{L}_{\text{IT-Align}}+\alpha\mathcal{L}_{\text{RC-Align}},$ (9)
where $\alpha$ is a hyper-parameter to adjust the weight of the RC-Align loss during pre-training.
@@
For zero-shot classification, our model integrates global and local alignment through dual-path scoring. Given an image $I_{m}$ with global representation $\mathbf{i}_{m,\text{cls}}\in\mathbb{R}^{h}$ and local region features $\mathbf{i}_{m}=[\mathbf{i}_{m,1},...,\mathbf{i}_{m,r}]\in\mathbb{R}^{r\times h}$, let class $c$ have global text embedding $\mathbf{t}_{c,\text{cls}}\in\mathbb{R}^{h}$ and $w_{c}$ concept embeddings $\mathbf{g}_{c}=[\mathbf{g}_{c,1},...,\mathbf{g}_{c,w_{c}}]\in\mathbb{R}^{w_{c}\times h}$.
Global scoring computes similarity between normalized global representations:
$s_{g}(c)=t_{g}\cdot\frac{{\mathbf{i}_{m,\text{cls}}^{\top}\mathbf{t}_{c,\text{cls}}}}{{\|\mathbf{i}_{m,\text{cls}}\|\|\mathbf{t}_{c,\text{cls}}\|}}+b_{g}.$ (10)
Local scoring aggregates region-concept information through pooled representations. Different from what we have done in the proposed RC-Align, we take a region selection criterion based on top-$k$ regions per concept:
$\mathbf{i}_{m,\text{loc}}$ $=\frac{1}{w_{c}k}\sum_{j=1}^{w_{c}}\sum_{i\in\mathcal{T}_{c,j}^{k}}\mathbf{i}_{m,i},\quad\mathbf{t}_{c,\text{loc}}=\frac{1}{w_{c}}\sum_{j=1}^{w_{c}}\mathbf{g}_{c,j},$ (11)
$s_{l}(c)$ $=t_{l}\cdot\frac{{\mathbf{i}_{m,\text{loc}}^{\top}\mathbf{t}_{c,\text{loc}}}}{{\|\mathbf{i}_{m,\text{loc}}\|\|\mathbf{t}_{c,\text{loc}}\|}}+b_{l},$
where the region selection criterion operates as follows:
$\forall i\in\mathcal{T}_{c,j}^{k},\forall n\notin\mathcal{T}_{c,j}^{k}:\frac{\mathbf{i}_{m,i}^{\top}\mathbf{g}_{c,j}}{\|\mathbf{i}_{m,i}\|\|\mathbf{g}_{c,j}\|}\geq\frac{\mathbf{i}_{m,n}^{\top}\mathbf{g}_{c,j}}{\|\mathbf{i}_{m,n}\|\|\mathbf{g}_{c,j}\|}.$
Here, $\mathcal{T}_{c,j}^{k}\subset\{1,\dots,r\}$ denotes the set of indices for the top-$k$ regions most similar to concept $j$ in class $c$.
Probability fusion combines both pathways:
$p_{g}(c)$ $=\frac{e^{s_{g}(c)}}{\sum_{d=1}^{C}e^{s_{g}(d)}},\quad p_{l}(c)=\frac{e^{s_{l}(c)}}{\sum_{d=1}^{C}e^{s_{l}(d)}},$ (12)
$p(c)$ $=\beta p_{l}(c)+(1-\beta)p_{g}(c).$
The final prediction $\hat{c}=\arg\max_{c}p(c)$ integrates both alignment strategies through learnable parameters $t_{g},b_{g},t_{l},b_{l}$ from Equations 3 and 6, with $\beta$ controlling the fusion ratio.
@@
In this section, we first describe the dataset used to pre-train the proposed ConceptCLIP model. Based on the pre-trained model, we conduct extensive experiments on downstream tasks. Besides, we describe the state-of-the-art medical multi-modal foundation models used for detailed comparison. Finally, we cover the implementation details in pre-training the proposed ConceptCLIP.
@@
In our pre-training process, the proposed MedConcept-23M is used. In detail, it contains 23 million image-text-concept triplets, where the image-text pairs are collected from PMC-OA and concepts are aligned to the UMLS concepts and extracted from the text.
Implementation details
We develop ConceptCLIP by utilizing SigLIP-ViT-400M-16111https://huggingface.co/timm/ViT-SO400M-14-SigLIP-384 as the visual encoder and PubMedBERT [^2] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as the text encoder. Each input image is resized to $336\times 336$ pixels. During pre-training, we employ the AdamW optimizer [33] with a learning rate of 0.0005. The model is initially pre-trained on 48 H800 GPUs with a batch size of 12,288 for 32 epochs on the MedConcept-23M dataset, without the RC-Align loss. Subsequently, we conduct further pre-training with the RC-Align loss on this dataset with a batch size of 6,144 and learning rate of 0.0003 for 20 epochs. The weight $\alpha$ (in Equation 9) during the pre-training is set to 0.5.
In medical image diagnosis, for zero-shot classification, we set $\beta$ (in Equation 12) to 0.5 so that the prediction can consider the information from both global information and local information. For linear probing, we use the LogisticRegression function provided in scikit-learn [^3] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, the inverse of regularization strength is 0.316, maximum iteration is 1,000, random state is 1, and all other parameters are default. For all fully fine-tuning experiments, the batch size is 64, epochs is 20, learning rate is 0.0001. We use AdamW optimizer for training the model. In the visual question answering task, we use AdamW as the optimizer. For SigLIP-400M and ConceptCLIP, the batch size is 32, learning rate is 0.000001. For other models, the batch size is 4, learning rate is 0.000005. All experiments have the same epoch number of 50. In pathology whole-slide image analysis, the batch size is 1, the learning rate is 0.0002, the epoch number is 30, and the cosine learning rate scheduler is used.
@@
Medical image diagnosis. We assess the zero-shot, linear probe (using 1%, 10%, and 100% of the training data), and fully fine-tuning (using 100% of the training data) capabilities of ConceptCLIP and previous state-of-the-art medical CLIP models across 36 datasets (Extended Data Table LABEL:tab:datasets) spanning 10 image modalities. In the zero-shot setting, no ground truth labels are provided. We follow the zero-shot protocol of the CLIP model [34], where the text encoder and image encoder are used to encode the medical images and labels (using a prompt template to form a sentence), respectively, and we then calculate the similarity and assign labels to the images. For linear probing, we first extract features from each image using a specific image encoder from a pre-trained model and employ these image features and their labels to perform logistic regression. The trained logistic regression models are subsequently used to test model performance. This experiment aims to evaluate each model’s ability to extract discriminative image features. In the fully fine-tuning setting, each model is concatenated with a classification head and fine-tuned on a specific dataset with all parameters adjustable. We use the AUC score for model evaluation to ensure that the performance is not influenced by thresholds.
@@
Cross-modal image-text retrieval. Cross-modal image-text retrieval includes tasks such as medical image-to-text and text-to-image retrieval. To evaluate the performance of ConceptCLIP across various medical image modalities, we utilize two datasets: PMC-9K and Quilt-1M [16], to assess the models’ cross-modal retrieval capabilities in a mix of image modalities and pathology images. We use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to evaluate each model’s performance.
@@
Medical report generation. Medical report generation aims to alleviate the workload of doctors by automatically generating a report given a medical image. In our study, following the shallow alignment in the R2GenGPT framework [35], we replace the vision encoder with that of each medical CLIP model and use the same training scheme to fine-tune the report generation model. We conduct these experiments on the MIMIC-CXR [17] and IU X-Ray [18] datasets. To comprehensively evaluate each model’s performance, we use natural language generation metrics: BLEU-1,2,3,4, METEOR, ROUGE-L, CIDEr, and clinical efficacy metrics: Micro Precision, Micro Recall, and Micro F1.
@@
Medical visual question answering. Medical Visual Question Answering (VQA) serves as a bridge between AI systems and humans. Given a medical image and a question about the image, the model provides an answer to the question based on the image. In our experiments, we use the METER framework [36] to evaluate different models on the VQA task, employing the SLAKE [37] and VQA-RAD [38] datasets. Following METER’s evaluation metrics, we use Closed Accuracy, Open Accuracy, and Overall Accuracy to assess each model’s performance.
@@
Pathology whole-slide image analysis. To explore our model’s ability to understand pathology whole-slide images, we conduct experiments on 6 related datasets for 9 subtasks in cancer diagnosis, mutation prediction, and survival prediction, each of which can be regarded as a classification task, with a whole-slide image as input to the vision encoder and a predicted label as output. Given the high resolution of whole-slide images, we follow the conventional approach by splitting the whole-slide image into smaller regions and using each pre-trained vision encoder to extract features from each image region. These features are input into the ABMIL model [39] for multiple instance learning (MIL). The model outputs a predicted label. Similar to the medical image diagnosis task, we use the AUC score to evaluate different models in cancer diagnosis and molecular subtyping, and the concordance index (C-index) for survival prediction.
@@
Medical concept annotation. We evaluate to what extent the proposed ConceptCLIP can annotate fine-grained concepts for medical images, hence facilitating the development of concept-based explainable artificial intelligence. The performance of medical concept annotation is evaluated using the concept labels of the datasets. For each concept, two positive and negative prompts are designed to calculate the similarity with input images to get the predicted output. Specifically, we adopt five datasets with clinical concept annotations to evaluate the zero-shot concept annotation performance, including dermatology images, blood cell images (pathology), breast ultrasound, and chest CT.
@@
Inherently interpretable model employment and evaluation. Building inherently interpretable models is essential for deploying AI in medicine since it is more trustworthy for doctors and patients during clinical diagnosis. In this paper, we employ the Concept Bottleneck Model (CBM) [9] to build the inherently interpretable model based on our method. Inspired by previous works [40, 41], we use ConceptCLIP to calculate the cosine similarity between input images and medical concepts, then the similarity is used as the input to the concept bottleneck layer, which maps the concept similarity to the final prediction. The concept bottleneck layer is trained using the class labels. For concept-class association, the weights of each neuron within the linear layer can be regarded as the contribution to the final prediction. To evaluate the effectiveness of our method, we compare ConceptCLIP with black-box models and other vision-language models of both the natural and medical domains on four datasets of different modalities, including dermatology, pathology, ultrasound, and CT.
Discovery for correlations between concepts and diseases. To conduct correlation discovery between concepts and diseases, concept presence difference is calculated for datasets of four modalities, including NLM-TB (chest X-ray), CBIS-DDSM (mammography), UBIBC (breast ultrasound), and MedFMC-Colon (pathology). Specifically, for each dataset, assume there are two image sets of different classes (e.g., tuberculosis and normal in NLM-TB dataset), denoted as $I_{+}=\{I_{1+},I_{2+},...,I_{M+}\}$ and $I_{-}=\{I_{1-},I_{2-},...,I_{N-}\}$, and the pre-defined concept list $C$ (e.g., cavitation, consolidation, etc.), where $M$ and $N$ are the numbers of images in the positive and negative image sets, respectively, and $C$ are generated by a large language model (i.e., GPT-4 [42]) since the dataset does not have concept labels. We adopt the proposed ConceptCLIP to annotate the presence of the concepts for each given image. The concept presence proportion of each image set is computed by $\frac{N_{c_{i}}}{M}$ and $\frac{N_{c_{i}}}{N}$, respectively, here $N_{c_{i}}$ is the number of images where a specific concept $c_{i}$ is present. Then the concept presence difference of concept $c_{i}$ is defined as the difference between the two concept presence proportions, i.e., $D_{c_{i}}=\frac{N_{c_{i}}}{M}-\frac{N_{c_{i}}}{N}$. A positive value of $D_{c_{i}}$ means concept $c_{i}$ is more likely to appear in the positive image set $I_{+}$, and vice versa. For disease-level concept inspection, as shown in Fig. 4 (i), pneumonia is selected as the inspected disease. For concept inspection on a dataset from one institution, the RSNA [43] dataset is adopted, which only includes classes of No finding and Pneumonia. For disease-level concept inspection, we collect datasets from various sources that also include the two classes, i.e., VinDr-CXR from the Hospital 108 and the Hanoi Medical University Hospital [44] and VinDr-PCXR from the Phu Tho Obstetric & Pediatric Hospital [45], where VinDr-PCXR focuses on the chest X-rays of children. It is noteworthy that VinDr-CXR and VinDr-PCXR contain other classes such as Lung tumor and Tuberculosis, but we only consider the images of No Finding and Pneumonia to analyze the pneumonia disease. For each dataset, we use ConceptCLIP to obtain the concept-disease association. Then we average the weights of the concept bottleneck layer across all considered datasets to derive the final top relevant concepts for pneumonia.
@@
AUC. The Area Under the Receiver Operating Characteristic Curve (AUC) is a performance measurement for classification problems at various threshold settings. It is widely used in medical image diagnosis and explainable AI tasks to evaluate the ability of a model to distinguish between classes.
Accuracy. Accuracy (ACC) quantifies the proportion of correct predictions relative to the total evaluated cases.
ACC=Number of Correct PredictionsTotal Number of $\text{ACC}=\frac{\text{Number of Correct Predictions}}{\text{Total Number of Predictions}}.$
For visual question answering (VQA) tasks, we evaluate three variants: closed accuracy, which measures performance on predefined answer candidates (e.g., multiple-choice questions); open accuracy, assessing free-form answer generation without predefined options; and overall accuracy, providing a combined assessment of both closed and open settings.
C-Index. The concordance index (C-Index) measures the predictive accuracy of a survival model. It is particularly used in survival prediction tasks in whole-slide image analysis to evaluate how well the model predicts the order of events. It is calculated as the proportion of all usable patient pairs whose predictions and outcomes are concordant.
Recall. Recall (sensitivity) quantifies the proportion of relevant instances successfully retrieved from the total relevant population. In classification tasks, this metric evaluates a model’s ability to minimize false negatives (FN), defined as:
$\text{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$